What would be the solution to the following system of equations:
2x + y - 3z = 11
-x + 2y + 4z = -3
x - 5y + 2z = -18
A.) (-1, 1, -4)
B.) (1, 3, -2)
C.) (3, 8, 1)
D.) (2, -3, 0)
Thank you in advance. :)

Answer by Edwin McCravy(20054)

2x +  y - 3z =  11
-x + 2y + 4z =  -3
 x - 5y + 2z = -18

Add the second and third equations 
to eliminate the x:

That will give you an equation with
only y and z in it

Then multiply the second original equation

-x + 2y + 4z =  -3
 
by 2 and add it the the first original equation

2x +  y - 3z =  11

and you will get another equation with
only y and z in it.  

[You will then be able to divide it through by 5.]

That will give you a system of 2 equations
equations in 2 unknowns, y and z.

Then you can solve that system of equations
and find y and z.  

Then finally you will substitute those values
for y and z into any one of the original three
equations and solve for x

If you have trouble, show what you've done in
the thank-you note form below and I'll get
back to you by email.

Edwin